OpenStudy (anonymous):

One pump can fill a tank with oil in 10 hours. Together with a second pump, it can fill the tank in 6 hours. How long will it take the second pump to fill the tank if it is used alone?

OpenStudy (mokeira):

in one hour, it will fill 1/10 how much will be filled if both work together? \[\frac{ 1 }{ 10 }+\frac{ 1 }{x }=\frac{ 1 }{6 }\] solve for x

OpenStudy (anonymous):

Add the rates.